Themodynamic Potentials Classical and Thermal Physics Helmholtz free energy dT-o applicationFr -pAtT HOMEWORK + A box consists of two chambers filled with(n and ng moles of an ideal gas separated by à movable impermeable wall. The box is kept at constant temperature inside a heat reservoir. gas gas reservoir
Potentials What is the relationship of the pressures on the two sides of the wal at equilibrium?V, dNo Find the location of the wall at equilibrium, i.e. what is L1/L2, the ratio of the widths of the chambers then?
Answer #1

1. At equilibrium, the membrane doesn't move, that means there shouldn't be any pressure difference between both sides of the membrane, so P- pressure on both sides of the membrane must be equal

2. ideal gas equation PV=nRT, since pressure in both chambers are equal and also since the box is maintained in constant temperature

V1/V2 = n1/n2 ; clearly V1/V2 =L1/L2

therefore L1/L2 = n1/n2